OpenStudy (anonymous):

rationalize the denominator (√(10)-√(3))/(√(10)+√(3))

OpenStudy (anonymous):

\[\frac{ \sqrt{10}-\sqrt{3} }{ \sqrt{10}+\sqrt{3} }\] This is the question

OpenStudy (mandre):

multiply by \[\frac{ \sqrt{10}-\sqrt{3} }{ \sqrt{10}-\sqrt{3} }\]

OpenStudy (anonymous):

@Mandre explain further mire please

OpenStudy (mandre):

The aim is just to rationalize the denominator. \[(\sqrt{10}+\sqrt{3})(\sqrt{10}-\sqrt{3}) = 10 - 3 = 7\] That would give you a rational denominator. Problem solved.

OpenStudy (mandre):

The numerator will be \[10-2\sqrt{10}\sqrt{3}+3\]